# Global Extraction of the $\rm{}^{12}C$ Nuclear Electromagnetic Response Functions (${\cal R}_{L}$ and ${\cal R}_{T}$) and Comparisons to Nuclear Theory and Neutrino/Electron Monte Carlo Generators

### v2

@@ V.1.2 Comparisons to ED-RMF in the nuclear excitation region
For example, the relative branching ratios for the decays of nuclear excitations between 12 and 21 MeV to protons and $\alpha$ particles have been measuredNeuschaefer and Tabor (1985). The decays of the 12.71 and 16.11 MeV excitations (which are below the proton and neutron separations energy) are primarily via the the emission of $\alpha$ particles, while for the 20.62 MeV excitation the branching ratio to decay via proton emission is a factor of 2 higher than the decay via $\alpha$ emission. Note that Christy-Bodek fit to inclusive electron scattering data describes all nuclear excitations irrespective of their decay modes.
@@ V.2.1 The NuWro Monte Carlo Generator
In the context of the ${}^{12}{\rm C}(e,e^{\prime})$ response functions, it is important to bear in mind that the electron mode of NuWro does not account for mechanisms of interaction other than QE scattering induced by the one-body current. The two-body currents and interference between one-body and 2-body currents are responsible for the enhancement of the transverse response functions. Therefore, in electron scattering mode NuWro is expected to underestimate the transverse response function in the QE region. and the comparisons with NuWro are only valid for ${\cal R}_{L}(\bf q,\nu)$
@@ V.2.2 Comparison to NuWro  predictions
The direct comparison with the measured values of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ as well as comparison with the Christy-Bodek universal fit can be used to extract corrections (e.g. tuning) that will bring the predictions into better agreement with the data as discussed in section 3.
@@ V.3 Comparison to the prediction of the Correlated Fermi Gas (CFG)
The recent Correlated Fermi Gas model Bhattacharya et al. (2024) includes two-nucleon correlations in a more sophisticated way. As seen in Fig.12 - 15 the model breaks down at low momentum transfers and works somewhat better at higher momentum transfers. Here also, the direct comparison with the measured values of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ as well as comparison with the Christy-Bodek universal fit can be used to extract corrections (e.g. tuning) that will bring the predictions into better agreement with the data as discussed in section 3.
@@ V.4.2 Comparison to achilles predictions
As seen in Fig.13 and Fig.14 the achilles  predictions for ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ are in better agreement with the data than the predictions of NuWro-SF-FSI. For ${\cal R}_{T}(\bf q,\nu)$ this is expected because achilles  also include the contribution of 2-body currents and the interference between 1-body and 2-body currents leading to 1-nucleon final states. The predictions for ${\cal R}_{T}(\bf q,\nu)$ are lower than the QE+TE component of Christy-Bodek fit because the transverse enhancement resulting in 2-nucleon final states (2p2h) are not included in the model.
@@
As expected, the NuWro predictions for ${\cal R}_{T}(\bf q,\nu)$ are underestimated because the electron mode of NuWro does not include the contribution of Meson Exchange Currents (which are included in the neutrino version of NuWro). Therefore, for NuWro the relevant validations are only for ${\cal R}_{L}(\bf q,\nu)$
@@ VI.2 SuSAV2 MEC-2p2h
In the Meson-Exchange-Current 2p2h model (MEC-2p2h) incorporated in SuSAv2 the main contribution in electron scattering is transverse and the contribution in the longitudinal channel is small. This is not the case for neutrinos where the 2p2h longitudinal axial contribution is important. In electron scattering the enhancement in the QE peak region originating from the 1p1h RMF transverse enhancement is in general bigger than the 2p2h contribution. The 2p2h channel becomes more important in the dip region between the QE and $\Delta$(1238) peaks. The sums of the predictions of SuSAv2-QE-1p1h and SuSAv2–MEC-2p2h are shown as the solid pink lines in the plots in Appendix A. In general the inclusion of the 2p2h contribution results in better agreement with our measurements in the dip region. However, at high values of q and $Q^{2}$ (e.g. $\bf q>$1.6 GeV and $Q^{2}>$ 1.25 GeV) the SuSAv2 MEC-2p2h model predictions for the longitudinal response functions have unphysical peaks. This could be because the parameterization of the 2p2h RFG-based nuclear responses from Ruiz-Simo et al.Ruiz Simo et al. (2017) included in this implementation of the SuSAv2-MEC model is valid only for ${\bf q}<2$ GeV and $\nu<2$ GeV (although it can be extended to higher kinematics).
In contrast the SuSAv2-MEC-2p2h implemented in genie  uses interpolation methods and therefore is more accurate than the parameterization formulae used in this implementation. This can introduce important differences at low kinematics (${\bf q}<$ 0.3 GeV) and also at very high kinematics (${\bf q}>$1.5 GeV).
@@
Model Currents Final State Validity RL RT high $\bf q$ low $\bf q$
ED-RMF 1b,2b 1p1h, nucl. exci. $all\leavevmode\nobreak\ \bf q$ OK OK 2p2h model 2p2h model
QMC no 2p2h need 2p2h need 2p2h needed needed
model model use SuSAV2 ? use SuSAv2?
SuSAV2 1b 1p1h + $all\leavevmode\nobreak\ \bf q$ low $\nu$ unphysical low $\nu$ unphysical OK need RL quench.
RMF Scaling for 2p2h model low q high low q low need more RT
function 1p1h no nucl exc. need quench
STA-QMC 1b,2b 1p1h, 2p2h $0.650.3<{\bf q}<$ 0.65 OK OK relativistic analytic extrapol.
no nucl exc. only corr. needed needed
NuWro SF-FSI 1b 1p1h all q too high too low needs needs
@@
achilles 1b,2b 1p1h ${\bf q}>$0.5 OK OK RL RT need other
QMC spectral no 2p2h. only need need need low q models
function no nucl exc. 2p2h model 2p2h model 2p2h model for 1p1h, 2p2h
GFMC 1b,2b 1p1h, 2p2h $0.570.3<{\bf q}<$ 0.57 low $\nu$ unphysical OK CPU intensive CPU intensive
QMC no nucl. exc. only q=0.57 too high not possible not possible
CFG 1b 1p1h,2p2h all q poor poor poor poor
@@
Because our extractions of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ from all available data electron scattering data span a large range of $\bf q$ and $Q^{2}$, the comparisons can be used extract $\bf q$ dependent corrections to nuclear models and neutrino/electron MC generators to achieve better agreement our measurement s (as has been done in the Christy-Bodek fit). For example, in the QE and nuclear excitation region:
@@
    The locations of the QE in $\nu$ are sensitive to the removal energy (energy shift) which shifts the QE peak to higher values of $\nu$) and also a $\bf q$ dependent final state optical potential Bodek and Cai (2019) which shifts the QE peak to lower values of $\nu$ at low $\bf q$. Tuning these parameters in SuSAv2 would bring the position of the QE peaks into better agreement with the measurements at low $\bf q$ and $Q^{2}$. In addition, introducing a q-dependence in the RMF-based SuSAv2 scaling functions (which are roughly constant for ${\bf q}>$ 0.3 MeV, but not for smaller values of q) can be used to incorporate into SuSAV2 the the low-energy RMF nuclear effects that break the scaling behavior at low q, which will modify the shape and magnitude of the SuSAv2 QE predictions at that region.
@@
    The location of the $\Delta$(1232) is also sensitive to removal energy and and final state optical potential. We observe that the shift of the peak of the $\Delta$(1232)  Bodek and Cai (2020) to lower values of $\nu$ is larger than the shift of the QE peak. This is unexpected because the optical potential for the $\Delta$(1232) is about 2/3 of the optical potential for the nucleonEricson and Weise (1988). In SuSAv2 this larger shift is accounted for by the 2p2h-MEC contribution to the cross section in the dip region. In gibuu this shift is accounted for by a nuclear pion production continuumMosel and Gallmeister (2024). In the Christy-Bodek fit this additional shift is parametrized in terms of a larger “effective optical potential” Bodek and Cai (2020) for the $\Delta$(1232) and inelastic pion production processes. Alternatively, the treatment of the real and imaginary parts of the $\Delta$ propagator in 2p2h and resonance channels and $\Delta$ in-medium modifications can also be tuned or modified to achieve better agreement with electron scattering data.
@@
    The magnitude of ${\cal R}_{L}(\bf q,\nu)$ can tuned be via a $\bf q$ dependent longitudinal quenching factor, and the amplitude of ${\cal R}_{T}(\bf q,\nu)$ can be tuned via an $\bf q$ dependent additive transverse enhancement contribution (as done in the Christy-Bodek fitBodek and Christy (2022)). Including and tuning such parameters in SuSAv2 would bring the magnitude of ${\cal R}_{L}^{QE}(\bf q,\nu)$ and ${\cal R}_{T}^{QE}(\bf q,\nu)$ predictions into better agreement with the measurements.
@@
  5. 5.
    A model describing nuclear excitations (with excitation energies below the proton separation energy) which decay via $\gamma$ ray emission is included in the inclusive Christy-Bodek fit Bodek and Christy (2023). At present none of the theoretical predications include nuclear excitations which decay via $\gamma$ emission.
@@ VII Summary
Since the extracted response functions cover a large range of $Q^{2}$ and $\nu$, they can be readily used to validate nuclear models as well as validate and tune Monte Carlo generators for electron and neutrino scattering experiments. Alternatively, the Christy-Bodek global fit to all electron scattering data on $\rm{}^{12}C$ can also be used to validate nuclear models and also validate and tune Monte Carlo generators for electron and neutrino scattering experiments over a larger region of $\nu$. In this paper we focus on comparison to models in the nuclear excitation region, and for single nucleon (QE-1p1h) and two nucleon (2p2h) final states.
@@
The STA-QMC calculation includes both single and two nucleon final states but at present it is only valid over a more restricted kinematic range and does not account for nuclear excitations. Extension of the theory to higher and lower values of momentum transfer are currently under way.
@@
VIII Supplemental Materials
The extracted ${\cal R}_{L}$ and ${\cal R}_{T}$ values (shown in the figures) are available as supplemental materials sup . In addition, the values of the Christy-Bodek fit to ${\cal R}_{L}$ and ${\cal R}_{T}$ over a larger range of $\nu$ (and also listing separately the contributions of nuclear excitations, quasielastic scattering, transverse enhancement, and pion production processes) are available as supplemental materials.
VIII.1 Acknowledgments

### v3

@@ V.1.2 Comparisons to ED-RMF in the nuclear excitation region
For example, the relative branching ratios for the decays of nuclear excitations between 12 and 21 MeV to protons and $\alpha$ particles have been measuredNeuschaefer and Tabor (1985). The decays of the 12.71 and 16.11 MeV excitations (which are below the proton and neutron separations energy) are primarily via the the emission of $\alpha$ particles, while for the 20.62 MeV excitation the branching ratio to decay via proton emission is a factor of 2 higher than the decay via $\alpha$ emission. Note that Christy-Bodek 2024 fit to inclusive electron scattering data describes all nuclear excitations irrespective of their decay modes.
@@ V.2.1 The NuWro Monte Carlo Generator
In the context of the $^{12}{\rm C}(e,e^{\prime})$ response functions, it is important to bear in mind that the electron mode of NuWro does not account for mechanisms of interaction other than QE scattering induced by the one-body current. The two-body currents and interference between one-body and 2-body currents are responsible for the enhancement of the transverse response functions. Therefore, in electron scattering mode NuWro is expected to underestimate the transverse response function and overestimate the longitudinal response function in the QE region
@@ V.2.2 Comparison to NuWro  predictions
The direct comparison with the measured values of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ as well as comparison with the Christy-Bodek 2024 universal fit can be used to extract corrections (e.g. tuning) that will bring the predictions into better agreement with the data as discussed in section 3.
@@ V.3 Comparison to the prediction of the Correlated Fermi Gas (CFG)
The recent Correlated Fermi Gas model Bhattacharya et al. (2024) includes two-nucleon correlations in a more sophisticated way. As seen in Fig.12 - 15 the model breaks down at low momentum transfers and works somewhat better at higher momentum transfers. Here also, the direct comparison with the measured values of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ as well as comparison with the Christy-Bodek 2024 universal fit can be used to extract corrections (e.g. tuning) that will bring the predictions into better agreement with the data as discussed in section 3.
@@ V.4.2 Comparison to achilles predictions
As seen in Fig.13 and Fig.14 the achilles  predictions for ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ are in better agreement with the data than the predictions of NuWro-SF-FSI. For ${\cal R}_{T}(\bf q,\nu)$ this is expected because achilles  also include the contribution of 2-body currents and the interference between 1-body and 2-body currents leading to 1-nucleon final states. The predictions for ${\cal R}_{T}(\bf q,\nu)$ are lower than the QE+TE component of Christy-Bodek 2024 fit because the transverse enhancement resulting in 2-nucleon final states (2p2h) are not included in the model.
@@
As expected, the NuWro predictions for ${\cal R}_{T}(\bf q,\nu)$ are underestimated because the electron mode of NuWro does not include the contribution of Meson Exchange Currents (which are included in the neutrino version of NuWro).
@@ VI.2 SuSAv2 MEC-2p2h
In the Meson-Exchange-Current 2p2h model (MEC-2p2h) incorporated in SuSAv2 the main contribution in electron scattering is transverse and the contribution in the longitudinal channel is small. This is not the case for neutrinos where the 2p2h longitudinal axial contribution is important. In electron scattering the enhancement in the QE peak region originating from the 1p1h RMF transverse enhancement is in general bigger than the 2p2h contribution. The 2p2h channel becomes more important in the dip region between the QE and $\Delta$(1238) peaks. The sums of the predictions of SuSAv2-QE-1p1h and SuSAv2–MEC-2p2h are shown as the solid pink lines in the plots in Appendix A. In general the inclusion of the 2p2h contribution results in better agreement with our measurements in the dip region. However, at high values of q and $Q^{2}$ (e.g. $\bf q>$1.6 GeV and $Q^{2}>$ 1.25 GeV2) the SuSAv2 MEC-2p2h model predictions for the longitudinal response functions have unphysical peaks. This could be because the parameterization of the 2p2h RFG-based nuclear responses from Ruiz-Simo et al.Ruiz Simo et al. (2017) included in this implementation of the SuSAv2-MEC model is valid only for ${\bf q}<2$ GeV and $\nu<2$ GeV (although it can be extended to higher kinematics).
Note that the SuSAv2-MEC-2p2h implemented in genie  uses interpolation methods and therefore is more accurate than the parameterization formulae used in this implementation. This can introduce important differences at low kinematics ($\bf q$}< 0.3 GeV) and also at very high kinematics ($\bf q$}>1.5 GeV).
VI.3 Resonance
@@
Model Currents Final State Available for RL RT Large $\bf q$ Small $\bf q$
ED-RMF 1b,2b 1p1h, nucl. exci. $all\leavevmode\nobreak\ \bf q$ best 1p1h best 1p1h best 1p1h best 1p1h
QMC 1b+2b best model 2p2h small need 2p2h need 2p2h need 2p2h
no 2p2h for 1p1h use SuSAv2 use SuSAv2 use SuSAv2 use SuSAv2
SuSAv2 1p1h 1p1h $all\leavevmode\nobreak\ \bf q$ low $\nu$ unphysical low $\nu$ unphysical OK needs RL
RMF Scaling only 1b only 1b 1p1h no 1b-2b low q model-high low q model-low quench
function 2p2h no nucl exc. interference needs 1p1h needs 1p1h needs RT
1b+2b 2p2h OK 2p2h OK quench or enhancement enhancement
use ED-RMF? use ED-RMF? use ED-RMF?
STA-QMC 1b,2b 1p1h + 2p2h. $0.3\leq{\bf q}<\leq.65$ OK OK relativistic analytic extrapol.
no nucl exc. $0.3\leq{\bf q}<\leq.65$ $0.3\leq{\bf q}\leq 0.65$ corr. needed needed
NuWro SF-FSI 1b 1p1h all q RL high RT low needs needs
@@
achilles 1b,2b 1p1h $\bf q$}>0.5 OK OK needs need other
QMC spectral no 2p2h. need need 2p2h model low q models
function no nucl exc. 2p2h model 2p2h model for RL RT for 1p1h, 2p2h
GFMC 1b,2b 1p1h+ 2p2h $0.3\leq{\bf q}\leq 0.57$ low $\nu$ unphysical OK CPU intensive CPU intensive
QMC no nucl. exc. q=0.57 high $0.3\leq{\bf q}<\leq.57$ not possible not possible
CFG 1b 1p1h+2p2h all q poor poor poor poor
@@
Because our extractions of ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ from all available data electron scattering data span a large range of $\bf q$ and $Q^{2}$, the comparisons can be used extract $\bf q$ dependent corrections to nuclear models and neutrino/electron MC generators to achieve better agreement our measurement s (as has been done in the Christy-Bodek 2024 fit). For example, in the QE and nuclear excitation region:
@@
    The locations of the QE in $\nu$ are sensitive to the removal energy (energy shift) which shifts the QE peak to higher values of $\nu$) and also a $\bf q$ dependent final state optical potential Bodek and Cai (2019) which shifts the QE peak to lower values of $\nu$ at low $\bf q$. Tuning these parameters in SuSAv2 would bring the position of the QE peaks into better agreement with the measurements at low $\bf q$ and $Q^{2}$. In addition, introducing a q-dependence in the RMF-based SuSAv2 scaling functions (which are roughly constant for $\bf q$}> 0.3 GeV, but not for smaller values of q) can be used to incorporate into SuSAv2 the the low-energy RMF nuclear effects that break the scaling behavior at low q, which will modify the shape and magnitude of the SuSAv2 QE predictions at that region.
@@
    The magnitude of ${\cal R}_{L}(\bf q,\nu)$ can tuned be via a $\bf q$ dependent longitudinal quenching factor, and the amplitude of ${\cal R}_{T}(\bf q,\nu)$ can be tuned via an $\bf q$ dependent additive transverse enhancement contribution (as done in the Christy-Bodek 2024 fitBodek and Christy (2022)). Including and tuning such parameters in SuSAv2 would bring the magnitude of ${\cal R}_{L}^{QE}(\bf q,\nu)$ and ${\cal R}_{T}^{QE}(\bf q,\nu)$ predictions into better agreement with the measurements.
@@
    A model describing nuclear excitations (with excitation energies below the proton separation energy) which decay via $\gamma$ ray emission is included in the inclusive Christy-Bodek 2024 fit Bodek and Christy (2023). At present none of the theoretical predications include nuclear excitations which decay via $\gamma$ emission.
@@ VII Summary
Since the extracted response functions cover a large range of $Q^{2}$ and $\nu$, they can be readily used to validate nuclear models as well as validate and tune Monte Carlo generators for electron and neutrino scattering experiments. Alternatively, the Christy-Bodek 2024 global fit to all electron scattering data on $\rm{}^{12}C$ can also be used to validate nuclear models and also validate and tune Monte Carlo generators for electron and neutrino scattering experiments over a larger region of $\nu$. In this paper we focus on comparison to models in the nuclear excitation region, and for single nucleon (QE-1p1h) and two nucleon (2p2h) final states.
@@
The STA-QMC calculation includes both single and two nucleon final states but is valid over a more restricted kinematic range and does not account for nuclear excitations. Extension of the theory to higher and lower values of momentum transfer are currently under way.
@@
The extracted ${\cal R}_{L}$ and ${\cal R}_{T}$ values (shown in the figures) are available as supplemental materials sup . In addition, the values of the Christy-Bodek 2024 fit to ${\cal R}_{L}$ and ${\cal R}_{T}$ over a larger range of $\nu$ (and also listing separately the contributions of nuclear excitations, quasielastic scattering, transverse enhancement, and pion production processes) are available as supplemental materials.
VIII Acknowledgments
@@ Appendix A Comparison to SuSAv2 in bins of $\bf q$ and $Q^{2}$
Comparisons to the predictions of SuSAv2 are shown in Figures 21-28.
Appendix B Quasi-deuteron contribution
The Quasi-deuteron contribution to the photoproduction ($Q^{2}=0$) cross section ($\sigma_{Quasi-D}$) as a function of $\nu=\varepsilon_{\gamma}$ is taken from Plujko et al. (2018). For a nucleus with $N$ neutrons, $Z$ protons and mass number $A=N+Z$, the $\sigma_{Quasi-D}$ is equal to
$\sigma_{Quasi-D}(\varepsilon_{\gamma})=397.8\;\frac{NZ}{A}\frac{(\varepsilon_{\gamma}-2.224)^{3/2}}{\varepsilon_{\gamma}^{3}}\phi(\varepsilon_{\gamma})$ (42)
with $\varepsilon_{\gamma}$ in MeV and $\sigma_{QD}$ in units of mb. The function $\phi(\varepsilon_{\gamma})$ accounts for the Pauli-blocking of the excited neutron-proton pair in the nuclear medium.
$\phi(\varepsilon_{\gamma}<20\;{\rm MeV})={\rm exp}(-73.3/\varepsilon_{\gamma}),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\$
$\phi(20<\varepsilon_{\gamma}<140\;{\rm MeV})=8.3714\times 10^{-2}$
$-9.8343\times 10^{-3}\varepsilon_{\gamma}+4.1222\times 10^{-4}\varepsilon_{\gamma}^{2}$
$-3.4762\times 10^{-6}\varepsilon_{\gamma}^{3}+9.3537\times 10^{-9}\varepsilon_{\gamma}^{4},$
$\phi(\varepsilon_{\gamma}>140\;{\rm MeV})={\rm exp}(-24.2/\varepsilon_{\gamma}).$ (43)
Appendix C Supplemental Materials
The following are available as supplemental materials sup
(a) Comparison of the normalized $^{12}C$ electron scattering cross sections to the universal fit (15 Ffgures).
(b) Tables of the extracted ${\cal R}_{L}$ and ${\cal R}_{T}$ values for bins of $\bf q$ and bins of $Q^{2}$.
(c) Tables of the values of universal fit values of ${\cal R}_{L}$ and ${\cal R}_{T}$ (the contributions of nuclear excitations, quasielastic scattering, transverse enhancement, and pion production processes are listed separately).